OpenStudy (anonymous):

Fan and Medal! Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3.

OpenStudy (freckles):

find g(-3) you are given g(x)=x-7

OpenStudy (freckles):

so replace x with -3 to get g(-3)

OpenStudy (anonymous):

g(-3) = -3 - 7 = -10

OpenStudy (freckles):

yes so now plug that number into f(x)

OpenStudy (anonymous):

f(-10) = 2(-10) + 5 = -20 + 5 = -15

OpenStudy (anonymous):

Thanks

OpenStudy (freckles):

np \[f(g(-3))=f(-3-7)=f(-10)=2(-10)+5=-20+5=-15 \] 

OpenStudy (anonymous):

Do you happen to know what solving this type of question is called?